OpenStudy (anonymous):

The sum of three integers is 136. The second number is three times the first. The sum of the first two numbers equals the third. Find the three integers.

OpenStudy (anonymous):

a + b + c = 136 3a = b a + b = c solve from there

OpenStudy (anonymous):

Yeah ive gotten that far i just dont know what to do from there..

OpenStudy (anonymous):

substitute 3a into a + b = c to get 4a = c then substitute for b & c with the equations you've got

OpenStudy (anonymous):

into the first

OpenStudy (anonymous):

Okay thanks!